# PolyFly: Polytopic Optimal Planning for Collision-Free Cable-Suspended Aerial Payload Transportation

### v1

@@ IV-C Distance Metric Between Polytopes
Our optimization problem seeks to lower bound $h_{M_{i},N_{j}}$, the minimum signed distance between polytopes $M_{i}$ and $N_{j}$, with a positive margin $\beta$. Specifically,
$h_{M_{i},N_{j}}(\mathbf{x})\geq\beta,~h_{M_{i},N_{j}}(\mathbf{x}):=sd(M_{i},N_{j}),$ (14)
where $sd$ is the signed distance function described in [19], which is positive when sets $M_{i}$ and $N_{j}$ do not intersect, and negative when they do. The constraint defined by eq. (14) is non-differentiable and difficult to integrate into optimization problems when the obstacles and robots are represented as convex polytopes [19]. To handle this complexity, we follow the approach in [19] and reformulate eq. (14) using duality theory into the following set of smooth non-linear differentiable constraints. The equivalent dual problem is
@@
$\mathbf{\lambda}_{M_{i}}\geq 0,\quad\mathbf{\lambda}_{N_{j}}\geq 0,\quad\left\|\mathbf{\lambda}_{M_{i}}*\mathbf{A}_{M_{i}}\right\|_{2}\leq 1$
Results in [19, 20] show that Strong Duality holds for the dual problem defined in eq. (15), which gives us the lower bound of the primal cost function $h_{M_{i},N_{j}}(\mathbf{x})$
@@
Therefore, to enforce the minimum signed distance of $\beta$ between $M_{i}$ and $N_{j}$, we constrain the lower bound $\bar{g}_{M,N}\geq\beta$, which gives us smooth non-conservative reformulation of eq. (14) as
@@
We add the set of constraints defined in eq. (17) for collisions between each $(M_{i},N_{j})$ pair, where $i\in(0,N_{O})$.
@@ IV-D Optimization Problem Formulation
$\min_{\mathbf{x}(k),\mathbf{u}(k),\Delta t(k)}$ $\sum_{k=0}^{N-1}L(\mathbf{x}(k),\mathbf{u}(k),\Delta t(k))$ (18)
subject to $\mathbf{x}(k+1)=F(\mathbf{x}(k),\mathbf{u}(k),\Delta t(k)),$
@@
$H_{l}(\mathbf{x}(k),\mathbf{u}(k))\leq 0,\forall l\in(0,N_{h}),$
where $N$ is the horizon length, $k$ is the stage index, $\Delta t(k)$ is the time duration for stage $k$, and $x_{0}$ and $x_{N}$ are the initial and terminal states respectively. $H_{l}$ comprises the $N_{h}$ inequality constraints from the collision-avoidance formulation in eq. (17) together with the state and input bounds specified in Section IV-D7. $L=L_{t}+L_{u}+L_{g}+L_{td}$ is the loss function defined later in this subsection. The optimization constraints, costs and initialization are discussed below.
@@ IV-D1 Time Minimization
$\begin{split}L_{t}&=\frac{\alpha_{to}}{N}\sum_{k=0}^{N-1}\Delta t_{k},\\ \Delta t_{min}&\leq\Delta t_{k}\leq\Delta t_{max},\end{split}$ (19)
@@ IV-D4 Regularizing Subsequent Time Step Durations
We add a small cost to the difference between subsequent time step durations to help our solver converge. Specifically, we add the following cost term
@@ IV-D5 Solver Initialization
Figure 2: A* initialization of trajectories for Env. 6 (left) and Env. 8 (right)
@@
$\dot{\mathbf{x}}_{L}{{}_{,k}}=\frac{\mathbf{x}_{L}{{}_{,k+1}}-\mathbf{x}_{L}{{}_{,k}}}{\Delta t_{max}-\Delta t_{min}}.$ (23)
Even with this relatively simple initialization method, our planner produces high quality trajectories that are consistently better than the baseline.
@@ IV-D6 Collision Constraints
We account for the collision between the quadrotor, cable, and payload polytopes against each obstacle. For each time stage $k$ and each polytope pair, we add the constraints derived in Section IV-C. As an example, to handle collisions between the payload polytope $N_{L}$ and an obstacles $M_{i}$, we add the following constraints for all stages $k$ and obstacles $i$
@@ IV-D7 State and Input Constraints
The initial and final positions are of the quadrotor and payload are specified based on the environment and are enforced via equality constraints
$\mathbf{x}(0)=\mathbf{x}_{0},\ \mathbf{x}(N)=\mathbf{x}_{N}.$ (25)
We constrain the start and end velocities, accelerations and jerks to zero to account for hardware limitations as
$\dot{\mathbf{x}}_{L}(0)=\dot{\mathbf{x}}_{L}(N)=0,~\ddot{\mathbf{x}}_{L}(0)=\ddot{\mathbf{x}}_{L}(N)$ $=0,$ (26)
$\dot{\mathbf{x}}_{Q}(0)=\dot{\mathbf{x}}_{Q}(N)=0,~\ddot{\mathbf{x}}_{Q}(0)=\ddot{\mathbf{x}}_{Q}(N)$ $=0,$ (27)
$\mathbf{u}(0)=\mathbf{u}(N)$ $=0.$ (28)
@@
$\mathbf{x}_{l}\leq\mathbf{x}\leq\mathbf{x}_{h},~\mathbf{u}_{l}\leq\mathbf{u}\leq\mathbf{u}_{h}.$ (29)
@@ IV-D8 Incorporating the Quadrotor’s Rotation into the Polytopic Constraints
To improve the accuracy of the polytopic representation, we rotate the quadrotor polytope by the quadrotor’s rotation. Due to the system’s differential flatness property, given $\mathbf{x}$ and $\mathbf{u}$, we can derive the robot’s rotation $\bm{\mathit{R}}_{Q}$ and use it to compute the $\mathbf{A}_{M_{i}}^{N_{Q}}$ matrix in eq. (13)
@@ V Experimental Results
Env. 1
Env. 2
Env. 3
Env. 4
Env. 5
Env. 6
Env. 7
Env. 8
Figure 3: Visualization of optimized trajectories using PolyFly in 8 tightly cluttered environments. The quadrotor is in yellow, payload in green, cable in gray and obstacles in light blue
We conduct various experiments to evaluate the capabilities of this method. We seek to answer the following questions:
@@
We create $8$ maze environments to evaluate the performance of our planning method. The optimization problem is constructed in CasADi [21] and solved it with IPOPT [22]. We compare our method against the MINCO approach [17] adopted in [6, 9] by utilizing their open-source code. To ensure a fair comparison, we adjust the baseline’s safety margins, ESDF map resolution, kinodynamic search resolution, and quadrotor parameters to be consistent with our environment setup. Furthermore, for each environment, we tune the baseline’s and PolyFly’s time minimization weights and report the best results in Table II. The optimized trajectories for our method in each environment are shown in Fig. 3.
The baseline method fails to produce a collision-free trajectory for $3/8$ environments, as indicated by the $\times$ symbol. We also observed that in some environments, the baseline generated shorter paths in terms of distance, yet required more time to traverse them. This indicates that our method is able to better optimize for total flight time, even if that entails following a slightly longer path at higher speeds. Across all test environments, PolyFly successfully generates paths where the robot tightly navigates around obstacles and aggressively cuts corners to maximize global performance. This boost in performance compared to the baseline partly stems from the geometric modeling of the workspace and the use of the dual formulation, which produces smooth, non-conservative collision constraints [19]. These properties enable the solver to plan paths that safely steer around obstacles while maintaining speed.
TABLE II: PolyFly vs. Baseline [6]. Time is total trajectory time and Path is the total path length. $\times$ marks a failure.
PolyFly Baseline[6] % Faster
Env. Time (s) Path (m) Time (s) Path (m)
1 3.92 4.46 4.10 4.45 4.1
2 5.02 7.17 $\times$ $\times$ $\infty$
3 4.95 6.65 5.68 6.47 12.3
4 7.17 8.81 7.47 8.74 4.3
5 4.20 5.74 4.30 4.92 6.3
6 5.45 8.08 6.43 8.02 15.7
7 3.16 3.05 $\times$ $\times$ $\infty$
8 4.93 9.42 $\times$ $\times$ $\infty$
@@
Modeling each robot component as a separate polytope allows the trajectory generator to accurately represent the system’s geometry. This fidelity is important when producing agile trajectories in tightly constrained environments, where the robot must cut corners aggressively while still avoiding obstacles. Moreover, the dual formulation of polytopic avoidance yields smooth, non-conservative constraints [19], in contrast to the baseline ESDF approaches, whose gradients can be non-smooth and discontinuous.
@@
The advantage is evident in Env. 7, where the robot must pass through a narrow opening wider than the cable but thinner than the payload and quadrotor. Our method successfully produces a trajectory that threads the cable through the gap, allowing the robot to safely navigate the environment. The real world experiment snapshots are shown in Fig. LABEL:fig:cover_page. The baseline fails to produce a feasible trajectory, likely due the geometric over-approximation and non-smooth gradients induced by the ESDF.
Q3 Incorporating the Quadrotor’s Rotation Into Polytopic Representation
@@
Accurately capturing the quadrotor’s rotation further increases the fidelity of the polytopic model and therefore navigation reliability. This detail becomes critical when the vehicle flies close to a ceiling, particularly for a platform with long rotor arms that forms a thin yet wide rectangular footprint. Since high-acceleration maneuvers induce large attitude changes, aggressive trajectories close to overhead surfaces can cause rotor collisions with the environment. As shown in Fig. 1, most planners approximate the vehicle as an ellipse, which is an over-approximation of the robot body, especially for large frames. By employing a tight polytope that explicitly accounts for the robot’s rotation, we eliminate this conservatism and ensure collision free navigation in near-ceiling flight.
@@
We demonstrate these benefits in a scenario where the quadrotor must fly just below a ceiling. Fig. 5 illustrates how incorporating the orientation into the planner prevents collisions. In Fig. 5 (A), the quadrotor strikes the ceiling during a high-acceleration maneuver because the planner ignores the attitude change that accompanies acceleration. In contrast, Fig. 5 (B) shows how the orientation-aware planner descends during the acceleration phase to avoid a collision.
Figure 4: Close up visualizations of the optimized trajectories through a narrow gap in Env. 7.
@@
Figure 5: Impact of incorporating quadrotor rotation in the polytopic model. Left: A rotation-agnostic planner produces a trajectory that collides with the ceiling. Right: By accounting for the robot’s attitude, the optimized path descends to prevent collisions while accelerating.
@@
We perform an ablation study to determine if regularizing the proximity to the initial guess is important. For each environment, we run the optimizer with $3$ different $\alpha_{g}$ values. The results are shown in Table III.
TABLE III: Trajectory time [s] for different values of $\alpha_{g}$. A $\times$ indicates a failure to generate a collision-free trajectory.
Env. $\bm{\alpha_{g}=0}$ $\bm{\alpha_{g}=1}$ $\bm{\alpha_{g}=5}$
1 $\times$ 3.92 3.92
2 $\times$ 5.01 5.02
3 $\times$ 3.94 4.95
4 $\times$ $\times$ 7.17
5 4.07 4.67 4.20
6 $\times$ 5.42 5.45
7 3.16 3.16 3.16
8 $\times$ 5.01 4.93
PolyFly fails in $6/8$ environments if $\alpha_{g}=0$ and in $1/8$ environments when $\alpha_{g}=1$. For $\alpha_{g}=5$, the optimizer succeeds in all environments. This highlights the value of this regularization term. We believe this aids convergence by incentivizing the optimizer to remain in the vicinity of the initial guess, preventing it from entering regions of local minima that are difficult to escape from.
@@
TABLE IV: Performance with and without velocity initialization across environments. A $\times$ indicates optimization failure.
@@
Env. Traj Time (s) Iterations Traj Time (s) Iterations
1 3.92 226 $\downarrow$ 3.93 418
2 5.02 411 $\downarrow$ 4.63 494
3 4.95 117 $\downarrow$ $\times$ $\times$
4 7.17 169 $\downarrow$ 7.14 756
5 4.20 402 $\downarrow$ 4.20 698
6 5.45 307 $\downarrow$ $\times$ $\times$
7 3.16 104 $\downarrow$ 3.18 126
8 4.93 192 $\downarrow$ 4.94 597
To evaluate the effectiveness of our solver initialization strategy, we conduct an ablation study that compares the payload-velocity initialization of eq. (23) against a variant where velocities are initialized to zero. Table IV summarizes the results.
The solver fails in $3/8$ environments without the velocity initialization strategy. For all other scenarios, our strategy on average significantly reduces the number of required solver iterations by roughly $45\%$.
@@
We run several real world experiments with environments $1$, 33 and $7$ to validate that our method produces trajectories that can be deployed on real robots. An overview of our system architecture is shown in Fig. 6. The desired $\mathbf{x}_{L}$, $\dot{\mathbf{x}}_{L}$, $\mathbf{x}_{Q}$ and $\dot{\mathbf{x}}_{Q}$ are sent to a Non-Linear Model Predictive Controller (NMPC) that runs onboard the robot at $200$ Hz. The NMPC uses the dynamics model presented in eqs. (1)–(5) with inputs $f_{1-4}$. We extract the NMPC’s expected thrust and bodyrates and send them to a low-level PX4 controller. The robot’s states are measured using an external motion capture system and are upsampled using onboard Kalman Filters.
Snapshots of our experiments are shown in Fig. LABEL:fig:cover_page. The position tracking errors for the quadrotor and payload are shown in Fig. 7 and Table V. Our results indicate that PolyFly produces dynamically feasible trajectories that can be tracked by an onboard controller with low errors. The supplementary video contains additional real-world footage that highlights the agile trajectories generated by our method.
Env. 1
Env. 3
Env. 7
Figure 7: Desired vs. Measured positions of the quadrotor and payload during real-world experiments for Env. 1, 3 and 7.
TABLE V: Tracking error statistics
Axis Error (m) Env. 1 Env. 3 Env. 7
$x$ Mean 0.027 0.029 0.021
Std 0.017 0.016 0.094
$y$ Mean 0.058 0.059 0.052
Std 0.047 0.048 0.036
$z$ Mean 0.024 0.039 0.029
Std 0.012 0.029 0.087

### v2

@@ IV-C Distance Metric Between Polytopes
Our optimization problem seeks to lower bound $h_{M_{i},N_{j}}$, the distance between polytopes $M_{i}$ and $N_{j}$, with a positive margin $\beta$. Specifically,
$h_{M_{i},N_{j}}(\mathbf{x})\geq\beta,~h_{M_{i},N_{j}}(\mathbf{x}):=d(M_{i},N_{j}),$ (11)
where $d$ is the distance function described in [25, Eq. (8a)]. The constraint defined by eq. (11) is non-differentiable and difficult to integrate into optimization problems when the obstacles and robots are represented as convex polytopes [25]. To handle this complexity, we follow the approach in [25, Proposition 1]. and reformulate eq. (11) using duality theory into the following set of smooth non-linear differentiable constraints. The equivalent dual problem is
@@
$\mathbf{\lambda}_{M_{i}}\geq 0,\quad\mathbf{\lambda}_{N_{j}}\geq 0,\quad\left\|\mathbf{\lambda}_{M_{i}}*\mathbf{A}_{M_{i}}\right\|_{2}\leq 1,$
where $\mathbf{\lambda}_{N_{j}}\in\mathbb{R}^{n_{j}}$ and $\mathbf{\lambda}_{M_{i}}\in\mathbb{R}^{m_{i}}$ are the dual variables associated with the obstacle avoidance problem. For our setup, $m_{i}=n_{j}=6$. Results in [25, 3] show that Strong Duality holds for the dual problem defined in eq. (12), which gives us the lower bound of the primal cost function $h_{M_{i},N_{j}}(\mathbf{x})$
@@
Therefore, to enforce a minimum distance of $\beta$ between $M_{i}$ and $N_{j}$, we constrain the lower bound $\bar{g}_{M,N}\geq\beta$, giving us a smooth non-conservative reformulation of eq. (11)
@@
We add the set of constraints defined in eq. (14) for collisions between each $(M_{i},N_{j})$ pair, where $i\in(0,N_{O})$. For notational convenience, we combine the dual variables into a vector $\mathbf{\Lambda}\in\mathbb{R}^{Q}$, where $Q=N\times N_{O}\times N_{R}\times(m_{i}+n_{j})$, and $N_{R}=3$ is the number of robot components.
@@ IV-D Optimization Problem Formulation
$\min_{\mathbf{x}(k),\mathbf{u}(k),\Delta t_{k},\mathbf{\Lambda}}$ $\sum_{k=0}^{N-1}L(\mathbf{x}(k),\mathbf{u}(k),\Delta t_{k})$ (15)
$\text{subject to \, \, \,}\mathbf{x}(k+1)$ $=F(\mathbf{x}(k),\mathbf{u}(k),\Delta t_{k},$
@@
$H_{l}(\mathbf{x}(k),\mathbf{u}(k),\mathbf{\Lambda})$ $\leq 0,~\Delta t_{min}\leq\Delta t_{k}\leq\Delta t_{max},$
where $\forall k\in(0,N-1)$, $\forall l\in(0,N_{h})$, $N$ is the horizon length, $\Delta t_{k}$ is the time duration for stage $k$, and $x_{0}$ and $x_{N}$ are the initial and terminal states respectively. $\Delta t_{min}$ and $\Delta t_{max}$ are the min. and max. time durations set by the user and are discussed in Section IV-D1. $H_{l}$ comprises the $N_{h}$ inequality constraints from the collision-avoidance formulation in eq. (14) together with the state and input bounds specified in Section IV-D7. Finally, $L=L_{t}+L_{u}+L_{g}+L_{td}$ is the loss function defined in the subsequent subsections.
We now describe the optimization constraints, costs and initialization methods. The variables $\Delta t_{min}$, $\Delta t_{max}$, $\alpha_{to}$, $\alpha_{u}$, $\alpha_{g}$, $\alpha_{td}$ and $N$, are user defined constants whose values are defined in Section V.
@@ IV-D1 Time Minimization
$L_{t}=T\frac{\alpha_{to}}{N},T=\sum_{k=0}^{N-1}\Delta t_{k},\Delta t_{min}\leq\Delta t_{k}\leq\Delta t_{max},$ (16)
@@ IV-D4 Regularizing Subsequent Time Step Durations
We add a small cost to the difference between subsequent time step durations to help our solver converge. Specifically, we add
@@ IV-D5 Solver Initialization
Figure 2: A* trajectories for Env. 6 (left) and Env. 8 (right)
@@
$\dot{\mathbf{x}}_{L}{{}_{,k}}=\frac{\mathbf{x}_{L}{{}_{,k+1}}-\mathbf{x}_{L}{{}_{,k}}}{\Delta t_{max}-\Delta t_{min}},$ (20)
where $0<\Delta t_{min}<\Delta t_{max}$. Even with this relatively simple initialization method, our planner produces high quality trajectories that are consistently better than the baseline.
@@ IV-D6 Collision Constraints
We account for the collision between the quadrotor, cable, and payload polytopes against each obstacle. For each time stage $k$ and each polytope pair, we add the constraints derived in Section IV-C. As an example, to handle collisions between the payload polytope $N_{L}$ and an obstacle $M_{i}$, we add the following constraints for all stages $k$ and obstacles $i$
@@ IV-D7 State and Input Constraints
The system’s initial and final positions are specified based on the environment and are enforced via equality constraints. We also constrain the start and end velocities, accelerations and jerks to zero to account for hardware limitations. Specifically,
$\mathbf{x}(0)=\mathbf{x}_{0},\ \mathbf{x}(N)=\mathbf{x}_{N},~\mathbf{u}(0)=\mathbf{u}(N)$ $=0,$ (22)
$\dot{\mathbf{x}}_{L}(0)=\dot{\mathbf{x}}_{L}(N)=0,~\ddot{\mathbf{x}}_{L}(0)=\ddot{\mathbf{x}}_{L}(N)$ $=0,$
$\dot{\mathbf{x}}_{Q}(0)=\dot{\mathbf{x}}_{Q}(N)=0,~\ddot{\mathbf{x}}_{Q}(0)=\ddot{\mathbf{x}}_{Q}(N)$ $=0.$
@@
$\mathbf{x}_{l}\leq\mathbf{x}\leq\mathbf{x}_{h},~\mathbf{u}_{l}\leq\mathbf{u}\leq\mathbf{u}_{h},$ (23)
where $\mathbf{x}_{l},\mathbf{u}_{l}$ and $\mathbf{x}_{h},\mathbf{u}_{h}$ are the lower and upper bounds.
@@ IV-D8 Incorporating the Quadrotor’s Rotation into the Polytopic Constraints
To improve the accuracy of the polytopic representation, we rotate the quadrotor polytope by the quadrotor’s rotation. As discussed in Section III, we can derive the robot’s rotation $\bm{\mathit{R}}_{Q}$ using $\mathbf{x}$ and $\mathbf{u}$, and use it to compute the $\mathbf{A}_{M_{i}}^{N_{Q}}$ matrix in eq. (10).
@@ V Experimental Results
Figure 3: Visualization of trajectories generated using PolyFly in 8 tightly cluttered environments. The quadrotor is in yellow, payload in green, cable in gray and obstacles in light blue. Envs. 1-4 are shown in the top row and 5-8 in the bottom row.
Figure 4: PolyFly’s trajectories for Env. 7 (top), 9 (bottom-left) and 10 (bottom-right) showcasing narrow-gap navigation.
We conduct various experiments to evaluate the capabilities of this method. We seek to answer the following:
@@
For all experiments, we set $\Delta t_{min}=0.01s$, $\Delta t_{max}=0.20s$, $\alpha_{u}=5.0$, $\alpha_{g}=5.0$, $\alpha_{td}=600.0$ and $N=100$. As described in Section V-Q1, $\alpha_{to}$ is tuned to maximize performance in each environment.
@@
We create $10$ maze environments to evaluate the performance of our planning method. The optimization problem is constructed in CasADi [1] and solved with IPOPT [19].
We select AutoTrans, the MINCO approach [21] adopted in [6, 20] as our primary baseline, which is an optimization-based trajectory generator that parametrizes trajectories as a sequence of $7^{th}$ order polynomials, initialized with kinodynamic hybrid $A^{*}$. They model the quadrotor and payload as independent spheres, the cable as a set of discrete points, and encode the environment using an ESDF. To solve the resulting problem, they reformulate the problem into an unconstrained one by applying $L_{1}$ penalties on the constraints.
To ensure a fair comparison, we adjusted AutoTrans’s safety margins, ESDF map resolution, kinodynamic search resolution, and quadrotor parameters to be consistent with our environment setup and removed AutoTrans’ hardware limitations constraints on the quadrotor thrust and tilt angle. Furthermore, for each environment, we tune AutoTrans’ and PolyFly’s time minimization weights and report the mean results based on the planning output over $10$ trials in Table II. PolyFly’s trajectories are shown in Fig. 3 and Fig. 4.
We include an additional baseline, PolySingle, which mirrors our method but adopts the modeling of [24], where a single polytope represents the entire robot. AutoTrans fails to produce a collision-free trajectory in $5/10$ environments, as indicated by a $\times$. We also observed that in some cases, AutoTrans generated shorter paths in terms of distance, yet required more time to traverse them. This indicates that our method is able to better optimize for total flight time, even if that entails following a slightly longer path at higher speeds. Compared to PolySingle, our method successfully navigates the narrow gaps in Env. 7, 9 and 10, whereas PolySingle fails due to its modeling over-approximation. PolyFly’s boost in performance compared to the AutoTrans may stem from the geometric modeling of the workspace and the use of the dual formulation, which produces smooth, non-conservative collision constraints [25]. These properties enable it to plan paths around obstacles while maintaining speed.
TABLE II: PolyFly vs. PolySingle vs AutoTrans. Statistics are calculated after running 10 trials in each environment. $T$ is the mean trajectory duration in seconds, Path is mean path length in meters, Comp. is the mean optimization wall time in seconds, and SR is the success ratio. $\%$ $T$ $\uparrow$ w/ AA denotes the $\%$ by which PolyFly’s trajectory duration $T$ is shorter than method AA’s. $\times$ marks a failure.
PolyFly PolySingle (PS) AutoTrans (AT)[6] % $T$ $\uparrow$ w/ AT % $T$ $\uparrow$ w/ PS
Env. $T$ Path Comp. SR $\bm{\alpha_{to}}$ $T$ Path Comp. SR $\bm{\alpha_{to}}$ $T$ Path Comp. SR
1 3.90 4.51 4.83 1.0 5000 3.96 4.50 5.26 1.0 5000 4.10 4.45 0.06 1.0 4.9 1.5
2 4.60 6.24 3.97 1.0 6000 4.68 6.34 2.99 1.0 6000 $\times$ $\times$ $\times$ 0.0 $\infty$ 1.7
3 4.92 6.24 7.11 1.0 5000 5.13 6.35 7.51 1.0 5000 5.68 6.47 0.11 1.0 13.4 4.1
4 7.09 8.82 6.80 1.0 6000 7.22 9.16 5.16 1.0 6000 7.47 8.74 0.89 0.9 5.1 1.8
5 4.05 5.82 11.14 1.0 5000 4.12 5.82 11.20 1.0 5000 4.30 4.92 1.63 0.6 5.8 1.7
6 5.31 7.95 4.90 1.0 6000 5.48 8.08 7.10 1.0 6000 6.43 8.02 49.44 0.0 17.4 3.1
7 2.86 3.07 3.12 1.0 800 $\times$ $\times$ $\times$ 0.0 800 $\times$ $\times$ $\times$ 0.0 $\infty$ $\infty$
8 4.92 9.40 3.31 1.0 10000 4.96 9.51 3.21 1.0 10000 $\times$ $\times$ $\times$ 0.0 $\infty$ 0.8
9 3.72 5.89 3.34 1.0 2000 $\times$ $\times$ $\times$ 0.0 2000 $\times$ $\times$ $\times$ 0.0 $\infty$ $\infty$
10 4.46 6.89 9.92 1.0 2000 $\times$ $\times$ $\times$ 0.0 2000 $\times$ $\times$ $\times$ 0.0 $\infty$ $\infty$
@@
Modeling each robot component as a separate polytope allows the trajectory generator to accurately represent the system’s geometry. This fidelity is important when producing agile trajectories to navigate through narrow gaps and in tightly constrained environments where the robot must cut corners aggressively while still avoiding obstacles.
@@
The advantage is evident in Env. 7, 9 and 10, where the robot must pass through a narrow opening wider than the cable but thinner than the payload and quadrotor. Both the baselines fail, but our method successfully produces trajectories that thread the cable through the gaps, allowing the robot to safely navigate the environment. The real world experiment snapshots are shown in Fig. LABEL:fig:cover_page.
PolySingle fails in these narrow-gap environments due to its single-polytope representation that over-approximates the robot’s geometry. On the other hand, AutoTrans models the cable as a discrete set of points and performs collision checks on each point, which should geometrically allow the cable to pass through the narrow gaps. Even though a solution exists, AutoTrans fails to produce a trajectory, even after extensive tuning of its optimization weights. Therefore, rather than the underlying geometry, the cause of the failures in AutoTrans may stem from other elements of the pipeline. First, the MINCO polynomial state parameterization, cannot, in general, represent optimal solutions [21] which limits its performance in tightly cluttered scenarios. And second, distance-field-based collision checking methods are not well-suited for applications that require accurate modeling of geometry [12]. These methods compute the gradients to prevent collisions per point of the robot/obstacle, rather than the exact gradient needed to move the polytope out of collision. These two gradients do not always agree, and can cause issues when the robot travels in close proximity to obstacles.
Q3 Incorporating the Quadrotor’s Orientation Into Polytopic Representation
@@
Accurately capturing the quadrotor’s orientation further increases the fidelity of the polytopic model and therefore navigation reliability. This detail becomes critical when the vehicle flies close to a ceiling or floor, particularly for a platform with long rotor arms that form a thin yet wide rectangular footprint. Since high-acceleration maneuvers induce large orientation changes, aggressive trajectories close to surfaces can cause rotor collisions with the environment. As shown in Fig. 1 and 6, most planners over-approximate the vehicle with a sphere or ellipsoid. By employing a tight polytope that accounts for the robot’s orientation, we eliminate this conservatism and ensure collision free navigation in near-ceiling flight.
@@
We demonstrate these benefits in Fig. 6 and 6. Fig. 6 illustrates how incorporating the orientation into the planner prevents collisions. In Fig. 6 (A), the quadrotor strikes the ceiling during a high-acceleration maneuver because the planner ignores the orientation change that accompanies acceleration. In contrast, Fig. 6 (B) shows how the orientation-aware planner descends during the acceleration phase to avoid a collision. The ceiling-floor scenario in Fig. 6 shows the PolyFly trajectory and overlays the PolySingle polytope to illustrate the over-approximation introduced by [24]. PolySingle fails in these scenarios because [24] uses a common orientation for all components due to its single polytope representation of the entire system, which prevents navigation in such tightly constrained environments.
@@
Figure 5: Left: An orientation-agnostic planner produces a trajectory that collides with the ceiling. Right: By accounting for the robot’s orientation, the optimized path descends to prevent collisions while accelerating.
Figure 6: Optimized trajectories by PolyFly for the ceiling-floor (left) and ceiling-wall (right) environments. The dashed lines are the polytopic approximation by PolySingle and the semi-transparent ellipse is the ellipsoidal approximation.
@@
We perform an ablation study to determine the importance of regularizing the proximity to the initial guess. We run the optimizer with $3$ different $\alpha_{g}$ values. The results are shown in Table III. PolyFly fails in $5/10$ environments if $\alpha_{g}=0$. Trajectories generated with $\alpha_{g}=1$ are on average slower than those generated with $\alpha_{g}=5$. This highlights the value of this regularization term. We believe this aids convergence by incentivizing the optimizer to remain in the vicinity of the initial guess, preventing it from entering regions of local minima that are difficult to escape from.
TABLE III: $T$ for different values of $\alpha_{g}$. % $T$ $\uparrow$ (5 w/ 1) indicates improvement of $\alpha_{g}=5$ over $\alpha_{g}=1$.
Env. $\alpha_{g}=0$ $\alpha_{g}=1$ $\alpha_{g}=5$ % $T$ $\uparrow$ (5 w/ 1)
1 $\times$ 3.92 3.92 0.0%
2 $\times$ 5.02 4.61 8.2%
3 4.94 4.96 4.97 -0.2%
4 $\times$ 7.13 7.15 0.3%
5 4.20 4.03 4.05 -0.5%
6 $\times$ 5.48 5.37 2.0%
7 3.15 4.59 3.16 31.2%
8 $\times$ 4.97 4.99 -0.4%
9 3.93 3.93 3.94 -0.3%
10 4.67 4.72 4.69 0.6%
@@
TABLE IV: Performance with and without $\dot{\mathbf{x}}_{L}$ initialization.
@@
Env. $T$ (s) Iterations $T$ (s) Iterations
1 3.90 201 $\downarrow$ 3.98 745
2 4.60 223 $\downarrow$ 4.61 640
3 4.92 133 $\downarrow$ 4.92 677
4 7.09 185 $\downarrow$ $\times$ $\times$
5 4.05 353 $\downarrow$ $\times$ $\times$
6 5.31 206 $\downarrow$ 5.31 559
7 2.86 234 $\downarrow$ 2.87 309
8 4.92 161 $\downarrow$ 4.90 617
9 3.72 112 $\downarrow$ 3.72 206
10 4.46 306 $\downarrow$ 4.43 652
To evaluate the effectiveness of our solver initialization strategy, we conduct an ablation study that compares the payload-velocity initialization of eq. (20) against a variant where velocities are initialized to zero. Table IV summarizes the results. The solver fails in $2/10$ environments without the velocity initialization strategy. For all other scenarios, our strategy reduces the solver iterations by an average of $59\%$.
@@
We run real world experiments with Env. $1$, 33 and $7$ to validate that our method produces trajectories that can be deployed on real robots. Our system architecture is shown in Fig. 7. For real-world experiments only, we use the Non-Linear Model Predictive Control (NMPC) setup from [10] as our controller. The NMPC uses the dynamics eqs. (1)–(5) with inputs $f_{1-4}$. After generating a trajectory with PolyFly, we send the desired $\mathbf{\tau}$ to the NMPC at $200$ Hz. As in [10], the predicted collective thrust and bodyrates from the MPC’s second stage are sent to the low-level controller. The robot’s states are measured using a motion capture system and are upsampled using Kalman Filters. Snapshots of our experiments are shown in Fig. LABEL:fig:cover_page. The position errors for the quadrotor and payload are shown in Fig. 8 and Table V. Our results indicate that PolyFly produces dynamically feasible trajectories that can be tracked by an onboard controller with low errors. The supplementary video contains real-world footage that highlights the agile trajectories generated by our method.
@@
Figure 8: Desired vs. Measured positions of the quadrotor and payload during real-world experiments for Env. 1, 3, and 7.
TABLE V: Tracking error statistics (meters)
Axis Env. 1 Env. 3 Env. 7
Mean Std Mean Std Mean Std
$x$ 0.027 0.017 0.029 0.016 0.021 0.094
$y$ 0.058 0.047 0.059 0.048 0.052 0.036
$z$ 0.024 0.012 0.039 0.029 0.029 0.087